OpenStudy (anonymous):

3 55/1000 please explain. rusty on my math

OpenStudy (barrycarter):

1000 has three zeros. When dividing by 1000, move the decimal to the left three places (add extra 0s if needed)

OpenStudy (anonymous):

the problem needs to be reduced

OpenStudy (anonymous):

If u meant 3(55/1000), i.e as a vulgar fraction (the fractions of the form p/q where p>q) it means 3+ 55/1000 = (3x1000 + 55)/ 1000 = 3055/ 1000 = 3.055. If u meant it as 355/1000, the answer will be 0.355 (i.e, move 3 places to left)